OpenStudy (anonymous):

What is the slope of the line represented by the equation 4x+3y=12 ?

OpenStudy (anonymous):

subtract 4x from both sides then divide by 3 on both sides, the number in front of x is the answer

OpenStudy (sweetnikkie072):

y=mx+b where m is slope 3y=12-4x y=(-4x+12)/3 y=(-4/3)x+4 therefore..-4/3 is ur slope

OpenStudy (anonymous):

slope is -4/3

OpenStudy (anonymous):



OpenStudy (anonymous):

are you understanding how to get the slope now?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

ok that's good